OpenStudy (alina123):

A scuba diver is 200 feet below sea level. Write a rational number to describe the divers depth. Part B in order to avoid decompression sickness , the diver must come up at a rate of 33/60 feet per second. How long will it take the diver to safely reach the surface? Show your work and round your answer to the nearest second.

OpenStudy (mathstudent55):

33/60 feet per second = 0.55 ft/sec speed = distance / time time = distance / speed distance = 200 ft speed = 0.55 ft/sec Find time.

OpenStudy (alina123):

Shouldnt the distance be -200 and im confused by what you wrote, can you explain it because when i divided the time didnt sound write to me

OpenStudy (mathstudent55):

Yes, -200 ft makes sense since it represents a distance below the surface of the sea.

OpenStudy (alina123):

but can you help me solve it

OpenStudy (alina123):

howd you get the speed by dividing distance

OpenStudy (mathstudent55):

Look at the equation for speed. To find a speed, you divide distance by time. s = d/t OK?

OpenStudy (alina123):

but how do you know the time

OpenStudy (alina123):

by divining the speed and distance

OpenStudy (mathstudent55):

Now take that equation, and solve it for time. s = d/t Multuply both sides by t: st = d Divide both sides be s: t = d/s Now you see that since speed equals distance divided by time, time equals distance divided by speed.

OpenStudy (mathstudent55):

Now divide the distance 200 ft by the speed 0.55 ft/sec to find the time.

OpenStudy (alina123):

i get 363.636363 so on

OpenStudy (mathstudent55):

Correct. That is in seconds. Now round off the answer as you asked to do.

OpenStudy (alina123):

round by which decimal place?

OpenStudy (alina123):

so 364 seconds